π 5. (14 pts.) Find the tangent plane when x =ī and y=- to the graph of z = 3 tan (2x - y).Find the tangent plane when and to the graph of.
Answer #1

Giver Z = 3tan (2x-y) 22 Zx= д д х (3 tan (21-у))3sec2(22-9)(2) 여로 (596sec2229) Z ( ) = 6sec/ 2(m) - )G (2 2 22 66V2) ? - 65ec²/ 741) 6(2)=12 +Zx (I, 7) = 12Zy - 7 dy بهای star (2x-y)) 2 (1-) (و-21)2 3Sec 2 (و ) و-27) 2 3re - Zy ( ره) -3 sec = ( 2 m) - )- 3 sec² (2(13)- ma -3sec 2/771) =-3 (12) 2 = -3/2) =-6 =-Zy (The II, 17-4)=-6 :: Equat on of tangent plane is:z-Z, = Zx(x,y) (x-x) + Zy(x,,,) (yg). Consider , Z = star (3x-y) 7 (I, I 3 tan ( 211 - TIS - 3tan = 31-1) =-3 Z, = -3Dㅕ +3) = 2z (1, 퓨)(x-1)+ u1 퓌 나퓌 12 (x-T) +(-)(-) +3 = 라 3= 121 -12T - 69f3 ㅠ 2.=). 12 x - by -Z 21T -3 so 2 Equation of targent plane is ZIT 12 X -64 - 7 - 3=0 2 12 1- 6Y - Z - 2117 3=0 2